What is the maximum wavelength of electromagnetic radiation that can eject photoelectrons from a metal with work function of 3.9 eV?


1
Expert's answer
2021-04-13T06:29:23-0400

The maximum wavelength of electromagnetic radiation that can eject photoelectrons from a metal is

"E_{phot}=\displaystyle \frac{hc} {\lambda_{max}}= A_{out}"

"\displaystyle \lambda_{max} = \frac{hc}{A_{out}}"

"\displaystyle \lambda_{max} = \frac{6.626 \cdot 10^{-34}\cdot 3 \cdot 10^8}{3.9 \cdot 1.6 \cdot 10^{-19}} = 3.185 \cdot 10^{-7} = 318.5 \cdot 10^{-9} \;m = 318.5 \; nm."


Answer: 318.5 nm.